[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Dec 07, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Thu Dec 7 12:58:44 UTC 2017


Here is today's summary of NHC computer operations:
--- Dave

----------

TAFB:
1. The ETSS header changed with an upgrade to the model yesterday. This did  
not affect NAWIPS, but prevented NHC and other sites from ingesting into  
AWIPS. NCF called, originally believing that the model was re-transmitted  
from NHC. Chris Mello was able to confirm that this is not the case. Chris  
made edits to NHC's AWIPS ingest to pull in the model but edex will need to  
be restarted at some point to commit the change. Since this is not  
operationally critical, that will be coordinated with the forecasters.

2. RWPS still missing- likely due to ongoing work on WCOSS. Regular  
activity on the dev machine resumed briefly yesterday but has been  
restricted again today for further testing. According to the latest  
schedule we have, regular access should resume this evening.
HSU:
None
Other items:
None
NHC RFCs:
None

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171207/6fffc836/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list